# ReDiFine: Reusable Diffusion Finetuning for Mitigating Degradation in the Chain of Diffusion

### v1

@@ Appendix D ReDiFine
Figure 25, 26, 27, and 28 show how robust ReDiFine is to different CFG scales. It successfully mitigates the high-frequency degradations for an extensive range of CFG scales.
@@
In addition to visual inspections for images generated using our ReDiFine, we also compare the quantitative results of ReDiFine to baselines with different CFG scales on FID, CLIP, and SFD. Results are shown in Figure 29. For different datasets and metrics, ReDiFine achieves comparable performance to the optimal CFG scale ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly).
@@
Appendix E Ablation study
@@
E.1 Condition drop finetuning
We conducted an ablation study to understand how the probability of dropping text embedding during finetuning affects the image quality in the Chain of Diffusion. Based on the probability used for training Stable Diffusion ($0.1$ or $0.2$), we tested $0.1$, $0.2$, and $0.4$. For both Pokemon and CelebA-1k datasets, a probability of 0.2 seems to work the best, as shown in Figure 30 and Figure 31, respectively. Interestingly, condition drop finetuning helps to mitigate the color saturation problem, but its effect decreases with a higher probability. For both of these datasets, condition drop finetuning can mitigate image degradation to some degree, but still, there is a large quality degradation that needs to be improved.
@@ E.1 Condition drop finetuning
E.2 CFG scheduling
We also tested how different decreasing strategies for CFG scale affect image degradation in the Chain of Diffusion. We used two different exponential orders for exponential decreasing, and also compared a linear decreasing strategy. Figure 32 shows that CFG scheduling is effective for Pokemon dataset, generating high-quality images comparable to those of ReDiFine. However, it fails to improve image quality on CelebA-1k dataset, as shown in Figure 33. This demonstrates the necessity of condition drop finetuning to improve the Chain of Diffusion universally for various datasets.
@@ E.2 CFG scheduling
Appendix F Analysis
In this section, we provide various analyses of images generated through the Chain of Diffusion. Specifically, we provide frequency analysis of latents during diffusion steps and fingerprints of synthetic images for forensic analysis.
F.1 Latent analysis
(a) Histogram of latent vectors.
(b) Diffs.
(c) Power spectrum density.
Figure 34: Fingerprints of the original training set and synthetic sets for Pokemon dataset. There are clear differences between fingerprints of natural images (Original) and synthetic images. ReDiFine makes similar fingerprints to CFG scale $2.5$. (a) Autocorrelation. Horizontal and vertical lines gradually disappear for CFG scales $1.0$ and $7.5$ while they are maintained for CFG scale $2.5$ and ReDiFine. (b) Average power spectra. Central regions are amplified or diminished for CFG scales $1.0$ and $7.5$, demonstrating low and high-frequency degradation.
Figure 34 demonstrates how latent distribution evolves for different iterations and diffusion steps on Pokemon dataset. First, the histogram of the final latent vectors before decoding is shown in Figure 3434(a). For a CFG of $1.0$, the latent distribution quickly evolves into a Gaussian-like shape, with its variance decreasing over iterations. This behavior aligns with previous work [5, 12, 19], which theoretically predicted that the self-consuming loop would crop out the tails of the distribution, reducing output diversity until it collapses to a single mode. We believe that these narrowing distributions in the latent space lead to blurrier and more homogeneous outputs in the image space. The evolution of the latent distribution is completely opposite for $7.5$, creating longer tails and approaching a more uniform distribution over space. For a CFG scale of $2.5$, which has the best reusability among the three, the latent distribution is better preserved over iterations. ReDiFine better preserves the histogram from the first to the last iteration, achieving both first iteration fidelity and reusability.
@@ F.1 Latent analysis
Figure 3434(c) demonstrates how the power spectrum density is maintained by ReDiFine. While high and low CFG scales cannot preserve the power spectrum of the first iteration as the Chain of Diffusion proceeds, ReDiFine can maintain the density distribution during iterations. This prevents images from being saturated to high or low-frequency degradation. Moreover, the power spectrum is more similar to the original training set compared to CFG scale $2.5$.
In summary, it is shown that ReDiFine can successfully alleviate the symptoms we observed in the latent space for baselines in Section 3. We can notice that ReDiFine preserves the original latent distribution at the $6$th iteration as compared to other baselines with different CFG values. More importantly, the conditional-unconditional score differences (Diff) are kept at bay all the way until the 6th iteration, as opposed to the baseline CFG $7.5$, showing a severe discrepancy between the unconditional and conditional scores.
F.2 Fingerprints for forensic analysis
@@
Fingerprints of synthetic images have distinct characteristics that differ according to the types of generative models, architectures, and training images. This provides unique patterns that can be used for synthetic image detection. Figure 9 shows how fingerprints of the original training set and synthetic images differ and how fingerprints of synthetic images change in the Chain of Diffusion. We use $1008$ images from Pokemon dataset. Following visual inspection, ReDiFine makes similar fingerprints to CFG scale $2.5$. Moreover, fingerprints for ReDiFine and baseline with CFG scale $2.5$ are maintained through the Chain of Diffusion. On the other hand, low or high CFG scales remain different patterns to the synthetic images as iterations proceed. Fingerprints gradually change for low or high CFG scales, showing the effects of high or low-frequency degradation.
Figure 35 demonstrates the radial and angular spectrum power density of the original and synthetic images. It is clearly observed that ReDiFine maintains a similar radial spectrum power density compared to the original training set, while even CFG scale $2.5$ fails. Moreover, it shows stable angular spectrums during the Chain of Diffusion despite it being distinct from that of the original training set.

### v2

@@ Appendix E ReDiFine
Figure 24, 25, 26, and 27 show how robust ReDiFine is to different CFG scales. It successfully mitigates the high-frequency degradations for a wide range of CFG scales.
@@
In addition to visual inspections for images generated using ReDiFine, we compare the quantitative results of ReDiFine to baselines with different CFG scales across FID and CLIP score. Furthermore, we compute the average sample-wise feature distance (SFD) between a pair of images corresponding to the text prompt as the fidelity metric applicable for text-to-image generation,
${\rm SFD}_{k}={\frac{1}{N}}\sum d(f(x_{k,i}),f(x_{0,i}))$ (17)
to evaluate each iteration $k$. SFD overcomes the problem of FID being sensitive to the number of images to compare. DiNOv2 features are used to compute SFD. The results, shown in Figure 28, demonstrate that ReDiFine achieves performance comparable to the optimal CFG scales ($2.5$ for Pokemon and Kumapi, $1.5$ for CelebA-1k and Butterfly) across different datasets and metrics.
@@
Appendix F Ablation study
@@
F.1 Condition drop finetuning
We conducted an ablation study to understand how the probability of dropping text embedding during finetuning affects the image quality in the Chain of Diffusion. We examine $0.1$, $0.2$, and $0.4$ as Stable Diffusion is trained using $0.1$ or $0.2$. For both Pokemon and CelebA-1k datasets, a probability of 0.2 works the best, as shown in Figure 29 and Figure 30, respectively. Interestingly, condition drop finetuning helps to mitigate the color saturation problem, but its effect decreases with a higher probability. For both of these datasets, condition drop finetuning can mitigate image degradation to some degree, but still, there is a large quality degradation that needs to be improved.
@@ F.1 Condition drop finetuning
F.2 CFG scheduling
We also evaluated how different CFG scale decreasing strategies impact image degradation in the Chain of Diffusion. We experimented with two different exponential decay rates and compared them with a linear decreasing strategy. Figure 31 demonstrates that CFG scheduling is effective for Pokemon dataset, generating high-quality images comparable to those generated by ReDiFine. However, as shown in Figure 32, it fails to enhance image quality on CelebA-1k dataset. This highlights the necessity of condition drop finetuning for achieving universal improvements in the Chain of Diffusion across various datasets.
@@ F.2 CFG scheduling
Appendix G Analysis
In this section, we present a series of analyses of images generated through the Chain of Diffusion. Specifically, we examine the distribution of latent values and the differences between conditional and unconditional scores. Additionally, we analyze the power spectra of the images using 2D Fourier transforms and explore fingerprints through forensic analysis [45, 46].
Figure 33: Histogram of latent values and Diffs during diffusion steps for Pokemon dataset. (a) Latent distribution shrinks over iteration for low CFG and expands with high CFG. Larger values in latent vectors are more likely to occur with high CFG, gradually increasing the tail of the distribution. (b) Differences between conditional and unconditional scores increase as the training set is more degraded. Especially, high differences in the later diffusion steps can be a cause of high-frequency degradation.
G.1 Latent analysis
Figure $3$(a) illustrates how the distribution of latent values evolves across different iterations. The histograms show the final latent vectors before decoding into pixel space, comparing various CFG scales and ReDiFine. For a CFG of $1.0$, the latent distribution rapidly converges into a Gaussian-like shape, with its variance shrinking over iterations. This behavior is consistent with previous work [3, 16, 11], which theoretically predicted that the self-consuming loop progressively trims the tails of the distribution, reducing output diversity until it collapses to a single mode. We hypothesize that this narrowing in the latent space leads to blurrier, more homogeneous outputs in pixel space. Conversely, at a CFG scale of $7.5$, the latent distribution develops longer tails and tends toward a more uniform spread across space. A CFG scale of $2.5$, which demonstrates the best reusability among the three, better preserves the latent distribution over iterations. ReDiFine further enhances this preservation, maintaining the histogram from the first to the last iteration, thus achieving both high fidelity in the first iteration and better reusability.
G.2 Differences between conditional and unconditional scores
@@
G.3 Power spectra of 2D Fourier transforms
Figure 34 demonstrates the radial and angular spectrum power density of both the original and synthetic images. It is evident that ReDiFine closely maintains the radial spectrum power density of the original training set, whereas even a CFG scale $2.5$ falls short. Additionally, ReDiFine demonstrates stable angular spectra throughout the Chain of Diffusion, even though they differ from those of the original training set. Pokemon dataset is used for power spectra analysis.
@@ G.3 Power spectra of 2D Fourier transforms
G.4 Fingerprints for forensic analysis
(a) Autocorrelation.
(b) Average power spectra.
Several works [45, 46] aim to identify fingerprints of synthetic images. High-quality synthetic images from different generative models have clearly distinct fingerprints, showing the potential to be used for synthetic image detection. We analyze fingerprints of synthetic images for different CFG scales and iterations, and compare them to fingerprints of the original training set. Both autocorrelation and average power spectra show clear differences between the original training set and synthetic images, as shown in Figure 35. Moreover, how the fingerprints of synthetic images evolve throughout the Chain of Diffusion differ for ReDiFine and different CFG scales. Specifically, fingerprints of synthetic images from ReDiFine are similar to those of images from CFG scale $2.5$, while other CFG scales ($1.0$ and $7.5$) make fingerprints different from the first iteration as iterations proceed. Horizontal and vertical lines in autocorrelation gradually disappear and central regions in power spectra vary for further iterations. Also, the varying central regions in power spectra imply that low frequency features increase and decrease for CFG scale $1.0$ and $7.5$, respectively, aligning with visual inspections. Generating images with fingerprints similar to those of the original real images can be an interesting future direction to reduce the degradation in the Chain of Diffusion. Pokemon dataset is used for fingerprint analysis.